OpenStudy (anonymous):

f(x) = 4x^2 - x - 3.. If f(x) = -3, what is x?

OpenStudy (anonymous):

Those were both incorrect.

OpenStudy (shamil98):

you never pointed out that i was solving the wrong problem LOL

OpenStudy (shamil98):

4x^2 - x - 3 = -3 4x^2 - x = 0 x(4x-1) = 0 x = 0 or 4x - 1 = 0 4x = 1 x = 1/4